# New Sorting Algorithm Wave Sort (W-Sort)

### v2

@@ Abstract.
Modern comparison sorts like quicksort suffer from performance inconsistencies due to suboptimal pivot selection, leading to $O(N^{2})$ worst-case complexity, while in-place merge sort variants face challenges with data movement overhead. We introduce Wave Sort, a novel in-place sorting algorithm that addresses these limitations through a dynamic pivot selection strategy. Wave Sort iteratively expands a sorted region and selects pivots from this growing sorted portion to partition adjacent unsorted data. This approach ensures robust pivot selection irrespective of dataset size, guarantees a logarithmic recursion stack depth, and enables efficient in-place sorting. Our analysis shows a worst-case comparison complexity bounded by $O(N(\log N)^{2})$ with a small constant factor. Experimental results demonstrate that Wave Sort requires significantly fewer comparisons than quicksort on average (approximately 24% less) and performs close to the theoretical minimum, while also incorporating adaptive techniques for efficient handling of presorted sequences. Wave Sort offers a compelling alternative for applications demanding consistent, predictable, and in-place sorting performance.
Sorting algorithms, In-place sorting algorithms, Analysis of algorithms
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Theory of computation Sorting and searching
1. Introduction
@@
The primary cause of inconsistent performance in modern quicksort implementations (Sedgewick and Wayne, 2011a) is lopsided partitions resulting from suboptimal pivot values. This can lead to a significant increase in both the number of comparisons and the depth of the recursion stack, degrading performance from an average of $O(NlogN)$ to a worst-case of $O(N^{2})$. Due to the computational cost associated with selecting a perfectly balanced pivot, current implementations often employ heuristic methods, such as ”median of three”(Sedgewick and Wayne, 2011a). While these methods attempt to choose a pivot from a representative sample, their efficacy can diminish as the dataset size increases because the sample size remains constant and does not scale accordingly. To ensure robust pivot selection irrespective of dataset size, the selection process would ideally need to be dynamic, considering an increasingly larger portion of the data. However, this approach typically increases computational overhead.
@@
Standard merge sort(Sedgewick and Wayne, 2011b) guarantees a comparison time of $O(NlogN)$ and logarithmic stack depth, but it requires additional linear space $O(N)$ and involves $O(NlogN)$ data movements. While in-place merge sort variants(Kim and Kutzner, 2003) eliminate the need for additional space, the complexity of comparison and element movement often increases with higher constant factors.(Huang and Langston, 1988)
@@
We introduce wave sort, a divide-and-conquer, in-place sorting algorithm designed to address the aforementioned shortcomings of quicksort and merge sort. The algorithm commences by sorting an initial small portion of the data. It then utilizes the median value of this sorted portion as the pivot to partition and sort progressively larger segments of the data until the entire dataset is sorted. This strategy effectively expands the algorithm’s ”frame of reference” for selecting a balanced pivot by leveraging already sorted sections of the array. Adaptive sorting is integrated to handle presorted sequence. It is also easy to effectively trade off comparison for element movement. While leapfrogging samplesort(Albacea, 1995) and some variations(Chen, 2007) employ a similar concept, their implementations often lack efficient block swap mechanisms, potentially leading to extra element moves. Furthermore, applying adaptive sorting techniques to these existing variations can be challenging.(Chen, 2006)
@@
Experimental work indicates that, on average, wave sort uses significantly fewer comparisons than quicksort (approximately 24% less) and a comparable number of element moves. The worst-case comparison complexity for wave sort is $O(N(logN)^{2})$, notably with a small constant factor (approximately $\frac{1}{4}$). Adaptive sorting techniques are integrated into wave sort to efficiently handle partially or fully presorted sequences. Additionally, the recursion stack depth is inherently restricted to logarithmic depth. Unlike introsort(Lammich, 2020), wave sort achieves this without requiring explicit monitoring of the stack depth.
2. Core Algorithm
@@
The algorithm is named wave sort because its active working set dynamically expands and contracts during execution, resembling the motion of a wave as it processes the data. Wave sort comprises two primary operations: up-wave and down-wave. The up-wave operation is responsible for expanding the currently sorted region to incorporate an adjacent unsorted portion of the data. Once this expanded working set (containing both the established sorted part and the newly included unsorted part) is defined, it is passed to the down-wave operation for sorting. The down-wave operation then sorts the unsorted portion within this working set using a divide-and-conquer approach. Crucially, pivots for partitioning the unsorted data are selected from the already sorted portion of the working set.
The overall process is iterative:
  1. (1)
@@
  • If the sorted portion (which acts as the pivot provider) is ”exhausted” (all its elements have been used as pivots, or its length is too short for effective partitioning) before the current unsorted portion is fully sorted, the remaining unsorted data is handled by initiating a new up-wave operation. This new operation will establish a new sorted base to continue sorting the remainder.
To simplify the description, and without loss of generality, the following assumptions are made:
  1. (1)
@@
  2. (2)
    The sorting process initiates from the rightmost end. An initial small segment at this end is sorted first, and then the ”wave” of sorting progresses towards the other end.
2.1. Driver function
This function serves as the initial entry point, handling the base case for sequences with fewer than two elements by simply returning. For sequences of two or more elements, it delegates the core sorting logic to the core algorithm.
1type seq []int
2
3func (s seq) WaveSort() {
4 if len(s) < 2 {
5 return
6 }
7 s.upwave(0, len(s)-1)
8}
@@ 2.1. Driver function
2.2. up-wave
Wave sort commences by establishing an initial sorted segment within the array. Analogous to dynamic programming, this establishes a base case or solves an initial ”subproblem.” This initial segment can be as small as a single element (which is inherently considered sorted) or a larger portion sorted using a conventional algorithm.
@@
Once the down-wave completes sorting this set, thereby enlarging the overall sorted region, the up-wave operation re-engages to expand the working set further. This cycle continues until the entire array is sorted. The up-wave thus serves as the main driver and entry point for the overall wave sort process. The mechanics of the up-wave process are illustrated in Fig. 1, and Listing 2.
1func (s seq) upwave(start, end int) {
2 if start == end {
3 return
4 }
5 sortedStart := end
6 sortedLength := 1
7 leftBound := end - 1
8 length := end - start + 1
9 for {
10 s.downwave(leftBound, sortedStart, end)
11 sortedStart = leftBound
12 sortedLength = end - sortedStart + 1
13 if length < sortedLength<<2 {
14 break
15 }
16 leftBound = end - sortedLength<<1 + 1
17 }
18 s.downwave(start, sortedStart, end)
19}
@@
Figure 1. up-wave
2.3. down-wave
Before diving into the details of the down-wave, it is necessary to introduce two sub-processes: partitioning and block swapping.
2.3.1. Partitioning
Partitioning splits the sequence into two parts: the left part contains elements less than the pivot, and the right part contains elements greater than the pivot. The function returns the index of the split point, which lies within the greater part.
Initialize pointer i to one position before the first element, and pointer j to one position after the last element. Let p be the pivot. Increment i until the element at index i is greater than or equal to the pivot. If i == j, return i as the middle index m. Then decrement j until the element at index j is less than or equal to the pivot. If j == i, return i as the middle index m. Swap the elements at indices i and j. Repeat this process. The partitioning process is illustrated in Fig. 2 and Listing  3.
1func (s seq) partition(l, r, p int) int {
2 i, j := l-1, r
3 for {
4 for {
5 i++
6 if i == j {
7 return i
8 }
9 if s.less(p, i) {
10 break
11 }
12 }
13 for {
14 j--
15 if j == i {
16 return i
17 }
18 if s.less(j, p) {
19 break
20 }
21 }
22 s.swap(i, j)
23 }
24}
@@ 2.3.1. Partitioning
Figure 2. Partitioning
2.3.2. Block swapping
Partitioning divides the unsorted portion into left and right blocks, separated by index m. The sorted portion is also split into left and right blocks, separated by the pivot. In the block swapping process, the right block of the unsorted portion (which becomes the left block in the swap) is exchanged with the left block of the sorted portion (which becomes the right block in the swap).
If the left block is shorter than the right block, define its length as $ll=r-m$, where r is the index of the first element in the right block, and m is the index of the first element in the left block. Set $nm=p-ll+1$, and initialize $j=m$. Store the element at index m in a temporary variable. Move the element at index $j+ll$ to index j, then update j to $j+ll$. If $j\geq nm$, replace the element at index $j+ll$ with the one at index $j-nm+m$. Repeat this process for $p-m+1$ iterations, as shown in Listing  4.
1func (s seq) blockSwap_sl(m, p, ll int) {
2 tmp := s[m]
3 init := m
4 j := m
5 nm := p - ll + 1
6 var k int
7 for range p - m + 1 {
8 if j >= nm {
9 k = j - nm + m
10 if k == init {
11 init++
12 s[j] = tmp
13 j = init
14 tmp = s[j]
15 continue
16 }
17 } else {
18 k = j + ll
19 }
20 s[j] = s[k]
21 j = k
22 }
23}
@@ 2.3.2. Block swapping
All together, the two scenarios are decided by the function shown in Listing  6.
1func (s seq) blockSwap(m, r, p int) {
2 ll := r - m
3 if ll == 0 {
4 return
5 }
6 lr := p - r + 1
7 if lr <= ll {
8 s.blockSwap_sr(m, r, p)
9 bs += lr << 1
10 return
11 }
12 s.blockSwap_sl(m, p, ll)
13}
@@
In both cases, each element is moved only once, with minimal use of a temporary variable. This ensures maximum efficiency. The block swap process is illustrated in Fig. 3.
Figure 3. block swapping
2.3.3. down-wave process
@@
Specifically, the elements to the left of $m$ in the unsorted portion, along with those to the left of the pivot in the sorted portion, belong to the left of the pivot in the final output. Conversely, the elements to the right of $m$ in the unsorted portion and to the right of the pivot in the sorted portion belong to the right of the pivot. A block swap is performed to exchange the right subset of the unsorted portion with the left subset of the sorted portion. This results in two new sections, each composed of an unsorted part on the left and a sorted part on the right.
The same partition and block swap operations are recursively applied to these new sections, progressively forming smaller subsections with the same structure. This continues until either the sorted or unsorted portion is fully exhausted. If the sorted portion is exhausted, an up-wave can be initiated to process the remaining unsorted portion, starting from its rightmost elements. If the unsorted portion is exhausted, the sorting process is complete. The down-wave procedure is illustrated in Fig. 4 and Listing  7.
1func (s seq) downwave(start, sortedStart, end int) {
2 if sortedStart-start == 0 {
3 return
4 }
5 p := sortedStart + (end-sortedStart)/2
6 m := s.partition(start, sortedStart, p)
7 if m == sortedStart {
8 if p == sortedStart {
9 s.upwave(start, sortedStart-1)
10 return
11 }
12 s.downwave(start, sortedStart, p-1)
13 return
14 }
15 s.blockSwap(m, sortedStart, p)
16 if m == start {
17 if p == sortedStart {
18 s.upwave(m+1, end)
19 return
20 }
21 p++
22 s.downwave(m+p-sortedStart, p, end)
23 return
24 }
25 if p == sortedStart {
26 s.upwave(start, m-1)
27 s.upwave(m+1, end)
28 return
29 }
30 s.downwave(start, m, m+p-sortedStart-1)
31 s.downwave(m+p-sortedStart+1, p+1, end)
32}
@@
Figure 4. down wave
3. Analysis
[…]
$p(i)=\frac{\binom{m}{i-1}\cdot\binom{m}{n-i}}{\binom{n}{2m+1}}$
[…]
$p(i)=\frac{\binom{m-1}{i-1}\cdot\binom{m}{n-i}}{\binom{n}{2m}}$
The probability of evenly dividing the unsorted portion is greater than $\frac{1}{n}$, as is the case in quicksort. Therefore, the average number of comparisons is expected to be closer to the best case than in quicksort. To simplify the analysis, assume the expansion ratio of the up-wave is 2; that is, each up-wave step doubles the size of the working set from the sorted portion. In the best case, the total number of comparisons for the largest up-wave working set is bounded by $C_{0}$:
$C_{0}=\frac{n}{2}+\left(\frac{n}{4}+\frac{n}{4}\right)+\cdots=\frac{n}{2}\log_{2}\frac{n}{2}$
The difference is in each division step of the down-wave, the pivot is removed from the sorted subset. The comparisons for the next smaller working set are bounded by $C_{1}$:
$C_{1}=\frac{n}{4}\log_{2}\frac{n}{4}$
Thus, the total comparison is less than $C_{b}$:
$C_{b}=C_{0}+C_{1}+...=nlog_{2}n-(n-1)$
In the worst case, each comparison provides the minimum information (lowest entropy). Assuming a sorted sequence for analysis, the total number of comparisons is denoted by $C_{w}$:
$C_{w}=\frac{n}{4}\left(\log_{2}^{2}n-\log_{2}\frac{n}{4}\right)$
The maximum stack depth is approximately $\log_{2}n$. In the worst case, the stack depth is less than:

### v3

# Wave Sort (W-Sort): A Novel In-Place Algorithm with Dynamic Pivoting for Consistent Performance and Logarithmic Stack Depth

@@ Abstract
Modern comparison sorts like quicksort suffer from performance inconsistencies due to suboptimal pivot selection, leading to $O(N^{2})$ worst-case complexity, while in-place merge sort variants face challenges with data movement overhead. We introduce Wave Sort, a novel in-place sorting algorithm that addresses these limitations through a dynamic pivot selection strategy. Wave Sort iteratively expands a sorted region and selects pivots from this growing sorted portion to partition adjacent unsorted data. This approach ensures robust pivot selection irrespective of dataset size, guarantees a logarithmic recursion stack depth, and enables efficient in-place sorting. Our analysis shows a best comparison complexity of N1N-1, average comparison complexity close to $\log_{2}(N)!$, and worst-case comparison complexity bounded by $O(N(\log(N))^{2})$ with a small constant factor, which could be reduced to $O(N\log(N))$ with hybrid sorting. The algorithm can be easily expanded to be hybridized with other sorting algorithms. Experimental results demonstrate that Wave Sort requires significantly fewer comparisons than quicksort on average (approximately 24% less) and performs close to the theoretical minimum $\log_{2}(N)!$. Wave Sort offers a compelling alternative for applications demanding consistent, predictable, and in-place sorting performance.
Introduction
@@
The primary cause of inconsistent performance in modern quicksort implementations (Sedgewick-quicksort) is lopsided partitions resulting from suboptimal pivot values. This can lead to a significant increase in both the number of comparisons and the depth of the recursion stack, degrading performance from an average of $O(N\log N)$ to a worst-case of $O(N^{2})$. Due to the computational cost associated with selecting a perfectly balanced pivot, current implementations often employ heuristic methods, such as the ”median of three” (Sedgewick-quicksort). Although these methods attempt to choose a pivot from a representative sample, their efficacy can diminish as the size of the data set increases because the sample size remains constant and does not scale accordingly. To ensure robust pivot selection, regardless of the size of the data set, the selection process would ideally need to be dynamic, considering an increasingly larger portion of the data. However, this approach typically increases computational overhead.
@@
The standard merge sort (Sedgewick-mergesort) guarantees a comparison time of $O(N\log(N))$ and logarithmic stack depth, but it requires additional linear space $O(N)$ and involves $O(N\log(N))$ data movements. Although in-place merge sort variants (in-place-mergesort) eliminate the need for additional space, the complexity of comparison and element movement often increases with higher constant factors (inplace-merging).
@@
We introduce wave sort, a divide-and-conquer, in-place sorting algorithm designed to address the aforementioned shortcomings of quicksort and merge sort. The algorithm begins by sorting an initial small portion of the data. Then it utilizes the median value of this sorted portion as the pivot to partition and sort progressively larger segments of the data until the entire dataset is sorted. This strategy effectively expands the algorithm’s ”frame of reference” for selecting a balanced pivot by leveraging already sorted sections of the array. The algorithm is divided into two logic components, up-wave and down-wave, to simplify the expansion, such as integration with adaptive sorting or hybrid with other sorting algorithms. Each component could have different expansion approaches. This paper shows that adaptive sorting is integrated to handle the pre-sorted sequence. Other adaptive sorting approaches, such as those shown in MBISort(MBISort), could also be integrated. It is also easy to effectively trade off comparison for element movement. Although leap-frogging samplesort (Leapfrogging-samplesort) and some variations such as partition sort (partition-sort) employ a similar concept, their implementations lack efficient block swap mechanisms, which could lead to extra element moves. Furthermore, applying adaptive sorting techniques to these existing variations can be challenging as shown in PEsort(PEsort).
@@
Experimental work indicates that, on average, wave sort uses significantly fewer comparisons than quicksort (approximately 24% less) and a comparable number of element moves. The worst-case comparison complexity for wave sorting is $O(N(\log(N))^{2})$, notably with a small constant factor (approximately $\frac{1}{4}$). Adaptive sorting techniques are integrated to efficiently handle partially or fully sorted sequences. Additionally, the recursion stack depth is inherently restricted to logarithmic depth. Wave sort could also replace quick sort in Introsort (introsort) to reduce the worst-case comparison complexity to $O(N\log(N))$. Wave sort makes the judgment of an unbalanced partition more accurate, so that it can avoid falsely falling into slower heapsort too early. The wave sort has $O(\log(N)^{2}$ stack depth in the worst case, so it is safe for Introsort to adapt more aggressive approaches.
1 Core Algorithm
@@
The algorithm is named wave sort because its active working set dynamically expands and contracts during execution, resembling the motion of a wave as it processes the data. The wave sort comprises two primary operations: up-wave and down-wave. The up-wave operation is responsible for expanding the currently sorted region to incorporate an adjacent unsorted portion of the data. Once this expanded working set (containing both the established sorted part and the newly included unsorted part) is defined, it is passed to the down-wave operation for sorting. The down-wave operation then sorts the unsorted portion within this working set using a divide-and-conquer approach. Crucially, pivots for partitioning the unsorted data are selected from the already sorted portion of the working set. The overall process is iterative:
  1. 1.
@@
  • If the sorted portion (which acts as the pivot provider) is ”exhausted” (all its elements have been used as pivots, or its length is too short for effective partitioning) before the current unsorted portion is fully sorted, the remaining unsorted data are handled by initiating a new up-wave operation. This new operation will establish a new sorted base to continue sorting the remaining.
To simplify the description and without loss of generality, the following assumptions are made:
  1. 1.
@@
  2. 2.
    The sorting process starts at the right end. An initial small segment at this end is sorted first, and then the ”wave” of sorting progresses towards the other end.
For brevity, we will be using the Go programming language programming:lang:go to illustrate the wave sort. The code complies with version 1.18 and above.
1.1 Driver function
This function serves as the initial entry point, handling the base case for sequences with fewer than two elements by simply returning. For sequences of two or more, it delegates the core sorting logic to the core algorithm.
type seq []int
func (s seq) WaveSort() {
if len(s) < 2 {
return
s.upwave(0, len(s)-1)
@@ 1.1 Driver function
1.2 up-wave
The wave sort commences by establishing an initial sorted segment within the array. Analogously to dynamic programming, this establishes a base case or solves an initial ”subproblem.” This initial segment can be as small as a single element (which is inherently considered sorted) or a larger portion sorted using a conventional algorithm.
@@
Once the down-wave completes sorting this set, thereby enlarging the overall sorted region, the up-wave operation re-engages to expand the working set further. This cycle continues until the entire array is sorted. The up-wave thus serves as the main driver and entry point for the overall wave-sorting process. The mechanics of the up-wave process are illustrated in Fig. 1, and Listing 2.
func (s seq) upwave(start, end int) {
if start == end {
return
sortedStart := end
sortedLength := 1
leftBound := end - 1
length := end - start + 1
for {
s.downwave(leftBound, sortedStart, end)
sortedStart = leftBound
sortedLength = end - sortedStart + 1
if length < sortedLength<<2 {
break
leftBound = end - sortedLength<<1 + 1
s.downwave(start, sortedStart, end)
@@
Figure 1: up-wave
1.3 down-wave
Before diving into the details of the down-wave, it is necessary to introduce two subprocesses: partitioning and block swapping.
1.3.1 Partitioning
Partitioning splits the sequence into two parts: The left part contains elements less than the pivot, and the right part contains elements greater than the pivot. The function returns the index of the split point that lies within the greater part.
Initialize the pointer i to the index of the position before the first element, and the pointer j to the index of the position after the last element. Let p be the pivot. Increment i until the element at index i is greater than or equal to the pivot. If i == j, return i as the middle index m. Then decrement j until the element at index j is less than the pivot. If j == i, return i as the middle index m. Then swap the elements at the indices i and j. Repeat this process. The partitioning process is illustrated in Fig. 2 and Listing 3.
The partitioning algorithm is different from the one from Sedgewick-quicksort and many other standard libraries that takes n+1 comparisons on an array of size n. It only takes n-1 comparisons on the same array and can be used to replace these algorithms to improve performance. The quicksort used to compare with the wave sort in this paper uses the same partitioning algorithm as the wave sort to eliminate unnecessary comparisons.
func (s seq) partition(l, r, p int) int {
i, j := l-1, r
for {
for {
i++
if i == j {
return i
if s.less(p, i) {
break
for {
j--
if j == i {
return i
if s.less(j, p) {
break
s.swap(i, j)
@@ 1.3.1 Partitioning
Figure 2: Partitioning
1.3.2 Block swap
Partitioning divides the unsorted portion into left and right blocks, separated by index m. The sorted portion is also divided into left and right blocks, separated by the pivot element. In the block-swapping process, the right block of the unsorted portion (which becomes the left block in the swap) is exchanged with the left block of the sorted portion (which becomes the right block in the swap).
If the left block is shorter than the right block, define the length of the left block as ll=rmll=r-m, where r is the index of the first element in the right block and m is the index of the first element in the left block. Set nm=pll+1nm=p-ll+1, and initialize $j=m$. Store the element at index m in a temporary variable. Move the element of index $j+ll$ to index j, then update j to $j+ll$. If $j\geq nm$, replace the element on the index $j+ll$ with the one on the index jnm+mj-nm+m. Repeat this process for pm+1p-m+1 iterations, as shown in Listing 4.
func (s seq) blockSwap_sl(m, p, $l$ int) {
tmp := s[m]
init := m
j := m
nm := p - $l$ + 1
var k int
for range p - m + 1 {
if j >= nm {
k = j - nm + m
if k == init {
init++
s[j] = tmp
j = init
tmp = s[j]
continue
} else {
k = j + $l$
s[j] = s[k]
j = k
@@ 1.3.2 Block swap
In this case, each element of the left and right blocks moves only once. The usage of the temporary variable is minimized. The total movement of the element is about the length of the left block ll plus the length of the right block lr. In a normal swap operation, there are 3 element moves. So each swap equals about 3 block swaps. In the case where the temporary variable could be squeezed into the register and the compiler could optimize the code to do so, there are two element moves in the swap. So each swap equals 2 block swaps in this case. In Leapfrogging-samplesort and its variation partition-sort to swap the left and right blocks, lr times swaps are required, which equals $2\times lr$ element moves or $3\times lr$ element moves if the register variable is not applicable. The block swap in this paper only needs $lr+ll$ element moves. For $ll<lr$, the moves of the element are reduced even if the register variable is applicable.
@@
In this case, the temporary variable is used only once. The moves of the element are the same as Leapfrogging-samplesort and its variation partition-sort if the register variable is applicable. Otherwise, the moves are reduced to $\frac{2}{3}$.
Together, the two scenarios are decided by the function shown in Listing 6.
func (s seq) blockSwap(m, r, p int) {
$l$ := r - m
if $l$ == 0 {
return
lr := p - r + 1
if lr <= $l$ {
s.blockSwap_sr(m, r, p)
return
s.blockSwap_sl(m, p, $l$)
@@
In both cases, each element concerned is moved only once, with minimal use of a temporary variable. Only the elements concerned are moved. This ensures maximum efficiency and eliminates the noise introduced in Leapfrogging-samplesort and its variation partition-sort for further analysis. The block swap process is illustrated in Fig. 3.
Figure 3: block swapping
1.3.3 Down-wave process
@@
Specifically, the elements to the left of $m$ in the unsorted portion, along with those to the left of the pivot in the sorted portion, belong to the left of the pivot in the final output. In contrast, the elements to the right of $m$ in the unsorted portion and to the right of the pivot in the sorted portion belong to the right of the pivot. A block swap is performed to exchange the right subset of the unsorted portion with the left subset of the sorted portion. This results in two new sections, each composed of an unsorted part on the left and a sorted part on the right.
The same partition and block swap operations are applied recursively to these new sections, progressively forming smaller subsections with the same structure. This continues until either the sorted or unsorted portion is fully exhausted. If the sorted portion is exhausted, an up-wave can be initiated to process the remaining unsorted portion, starting from its rightmost elements. If the unsorted portion is exhausted, the sorting process is complete. The down-wave procedure is illustrated in Fig. 4 and in Listing 7.
func (s seq) downwave(start, sortedStart, end int) {
if sortedStart-start == 0 {
return
p := sortedStart + (end-sortedStart)/2
m := s.partition(start, sortedStart, p)
if m == sortedStart {
if p == sortedStart {
s.upwave(start, sortedStart-1)
return
s.downwave(start, sortedStart, p-1)
return
s.blockSwap(m, sortedStart, p)
if m == start {
if p == sortedStart {
s.upwave(m+1, end)
return
p++
s.downwave(m+p-sortedStart, p, end)
return
if p == sortedStart {
s.upwave(start, m-1)
s.upwave(m+1, end)
return
s.downwave(start, m, m+p-sortedStart-1)
s.downwave(m+p-sortedStart+1, p+1, end)
@@
Figure 4: down wave
2 Analysis
Let the length of the sequence be $n$, and let the length of the sorted portion be $\mathit{ls}$. If $\mathit{ls}$ is odd, it can be expressed as $2m+1$. The probability that the median of the sorted portion appears at position $i$ in the sequence is:
$p(i)=\frac{\binom{i-1}{m}\cdot\binom{n-i}{m}}{\binom{n}{2m+1}}$ (1)
If $\mathit{ls}$ is even, it can be expressed as $2m$. In this case, the probability that the median appears at position $i$ is:
$p(i)=\frac{\binom{i-1}{m-1}\cdot\binom{n-i}{m}}{\binom{n}{2m}}$ (2)
The probability of evenly dividing the unsorted portion is greater than $\frac{1}{n}$, as is the case in quick sort. Therefore, the average number of comparisons is expected to be closer to the best case than in quicksort. To simplify the analysis, assume the expansion ratio of the up-wave is always 2; that is, each up-wave step doubles the size of the working set from the sorted portion. In the best case, the total number of comparisons for the largest up-wave working set is bounded by $C_{0}$:
$C_{0}=\frac{n}{2}+\left(\frac{n}{4}+\frac{n}{4}\right)+\cdots-E_{0}=\frac{n}{2}\log_{2}\frac{n}{2}-E_{0}$
The difference $E_{0}$ is that in each division step of the down wave, the pivot is removed from the sorted subset. So some of the branches may end earlier and the comparison only needs to be done among about half of the pairs that are approximately equal to $\frac{n}{8}$ or a little more. The comparisons for the next smaller working set are bounded by $C_{1}$:
$C_{1}=\frac{n}{4}\log_{2}\frac{n}{4}-E_{1}$
where $E_{1}=E_{0}/2\approx\frac{n}{16}$. Thus, the total comparison is $C_{b}$:
$C_{b}=C_{0}+C_{1}+...C_{m}-E=nlog_{2}n-(n-1)-E$
where $m=log_{2}\frac{n}{2}-1$ , $E=\sum_{i=0}^{m-1}E_{i}\approx\frac{n}{4}$.
$C_{b}\approx nlog_{2}n-(n-1)-\frac{n}{4}$
$=nlog_{2}n-\frac{5n}{4}+1$
In the worst case, each comparison provides the minimum information (lowest entropy). A sorted sequence fulfills this situation, so a sorted sequence is used for analysis and testing. The total number of comparisons is denoted by $C_{w}$. In this case, the number of comparisons for the largest up-wave working set is $\frac{n}{2}\log_{2}\frac{n}{2}$ and there is only one such working set. The comparisons for the next smaller working set are $\frac{n}{4}\log_{2}\frac{n}{4}$ and the number of such working sets is 2. So we get:
$C_{w}=1\times\frac{n}{2}\log_{2}\frac{n}{2}+2\times\frac{n}{4}\log_{2}\frac{n}{4}+4\times\frac{n}{8}\log_{2}\frac{n}{8}$
$+\cdots+\frac{n}{8}\times 4\log_{2}4+\frac{n}{4}\times 2\log_{2}2$
$=\frac{n}{4}\left(\log_{2}^{2}n-\log_{2}\frac{n}{4}\right)$
The maximum stack depth is approximately $\log_{2}n$ as in the quick sort. In the worst case, the stack depth is less than: